OpenStudy (anonymous):

PLEASE HELP ME!!! Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). (5 points)

OpenStudy (anonymous):

i know much how time but idk two linear equations!! @AriPotta @Catlover5925 @campbell_st

OpenStudy (anonymous):

35 mintues breast stroke and 15 minutes butterfly

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

@Secret-Ninja @KamiBug @Catlover5925

OpenStudy (anonymous):

@phi

OpenStudy (phi):

they tell you number of minutes Ted practices the butterfly stroke every day (x) in other words, "x" is the name for number of minutes swimming the butterfly also, y is the number of minutes swimming the breast stroke.

OpenStudy (phi):

if you add them up and use the info for a total of 50 minutes every day. can you write an equation ?

OpenStudy (anonymous):

but then the next part is how long does Ted swing the Breast stroke well its 35 and butterfly is 15 so y=15+35?

OpenStudy (phi):

First, can you write an equation for how much the sum of x and y is ?

OpenStudy (anonymous):

50=x+y

OpenStudy (phi):

think: "number of minutes on breast" plus "number of minutes on butterfly" is 50 yes 50= x+y or x+y= 50 (this is standard form... the variables are on the left side)

OpenStudy (anonymous):

okay so but it says two pairs of linear equations

OpenStudy (anonymous):

so x+y=50 and what is another one

OpenStudy (phi):

now 20 minutes more practicing the breast stroke than the butterfly now use common sense, which does he practice more on, breast or butterfly ?

OpenStudy (anonymous):

breast.

OpenStudy (phi):

butter fly is x breast is y y is more than x (more time on the breast) y is how many more minutes than x ?

OpenStudy (anonymous):

20

OpenStudy (phi):

so say to your self y is 20 more than x write that as an equation

OpenStudy (phi):

is --> = more than --> add

OpenStudy (anonymous):

20y+x=50?? but i dont think thats it

OpenStudy (phi):

y is 20 more than x you can usually think of "is" as = so, getting closer to an equation: y = 20 more than x

OpenStudy (anonymous):

i truthfully dont know

OpenStudy (phi):

what is one more than two ?

OpenStudy (anonymous):

3

OpenStudy (phi):

how did you figure that out ?

OpenStudy (anonymous):

because if you add 1 to two you get 3 lol

OpenStudy (anonymous):

50-20=x+y?

OpenStudy (phi):

yes. now use the same idea: one more than x

OpenStudy (phi):

you can't get a number, but you can show one more than x as 1+x

OpenStudy (anonymous):

okay

OpenStudy (phi):

it's like one more than two is 1+2 (which you can simplify to 3) so far we have y = 20 more than x any ideas?

OpenStudy (anonymous):

so y=x+20?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thanks!!!

OpenStudy (phi):

notice that y=x+20 is really a (very) shortened version of y (time spent on breast) is 20 minutes more than x (time spent on butterfly)

OpenStudy (phi):

to solve x+y= 50 y= x+20 you can use substitution. y is the "same" as x+20 so in the first equation, write it as x+ (x+20) = 50 and simplify... and solve for x